The probability of a cooker having a fault in the first year is 0.12.
a) what is the probability of a cooker not having in the first year? .
b) if the 2000 cookers are produced , how many would you expect to devolop a fault in the first year?
1
Expert's answer
2011-06-06T07:15:29-0400
The probability of a cooker having a fault in the first year is 0.12.
& a) What is the probability of a cooker not having in the first year?
Solution:
The probability of not having a fault is: (1 -& the probability of having a fault) = 1 - 0.12 = 0.08
Answer: 0.08
& b) If the 2000 cookers are produced, how many would you expect to develop a fault in the first year?
Solution:
Expected value = x*p(x) = 2000*0.12 = 240
Answer: 240